# Robotic Learning for Informative Path Planning

### v1

@@ 4.2 Reinforcement Learning
The problem of AIPP can be framed as a partially observable Markov decision process (POMDP)  (Kochenderfer et al., 2022). Formally, a POMDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\rangle$, where:
@@
In many real-world scenarios, the agent cannot directly observe the true state of the environment. Instead, it receives observations that provide partial information about the true state. In the context of AIPP, the true state of the environment is the actual physical attributes of the environment captured by $F(\mathbf{x})$ in Sec. 2, e.g., the locations of obstacles and areas of interest, while the observation $z$ is the robot’s current measurement of these attributes based on sensor readings.
@@
The state transition function $T(s^{\prime}\,|\,s,a)$ describes the probabilistic dynamics of the robot’s state in the context of the AIPP problem. When the robot executes an action $a\in\mathcal{A}$ in its current state $s\in\mathcal{S}$, it moves to a new state $s^{\prime}\in\mathcal{S}$. The inherent uncertainty of action outcomes and their impact on the environment is captured by the transition function. This function is used in the belief update, which refines the robot’s understanding of the environment feature mapping $F$. Additionally, the observation function $Z(o\,|\,s^{\prime},a)$ provides the likelihood of perceiving an observation $o\in\mathcal{O}$ subsequent to action $a$ and arriving at state $s^{\prime}$. This observation model, which is probabilistically linked to the feature mapping $F$ through $z_{t}\sim p(z\,|\,F)$, is used to update the robot’s belief state using Eq. 2. After executing action $a$ and observing $o$, the belief over the feature mapping $F$ is updated. This update enables the robot to better predict environmental features, which in turn aids in refining its path planning.
@@
Any POMDP can be viewed as an MDP that uses beliefs as states, also called a belief-state MDP or belief MDP. The state space of a belief MDP is the set of all beliefs $\mathcal{B}$. The action space is equivalent to that of the POMDP (Kochenderfer et al., 2022). To connect this directly to Sec. 3, in AIPP, the robot’s belief about the true state of the environment is denoted by $\tilde{F}$ in Sec. 3.1. This belief evolves as the robot gathers information about its environment, which starts as partially unknown and is gradually discovered through sensing and exploration. As discussed in Sec. 3.1, the structure of this belief involves the robot creating an environment map from the sensor measurements. The robot’s beliefs about the feature mappings, represented by $\tilde{F}$, are updated as it collects measurements and takes actions. These beliefs are then used to quantify the change in information criterion $I(\cdot)$. The performance of a specific policy can be evaluated using the metrics described in Sec. 3.2 based on the posterior map belief $\tilde{F}$.
@@
All of the methods surveyed include some aspect of the current robot information in the state. The robot pose is often used (Cao et al., 2023b; Lodel et al., 2022; Zeng et al., 2022; Marchant et al., 2014; Denniston et al., 2023b), while other methods include additional robot information, e.g., energy level (Choudhury et al., 2020; Ott et al., 2023), viewing direction (Chen et al., 2022), and operational status (Bayerlein et al., 2021). In addition to current robot information, all methods store some form of a spatial map in the state. The most popular spatial maps are Gaussian processes (Cao et al., 2023a, c; Yang et al., 2023a; Wei and Zheng, 2020; Rückin et al., 2022a; Ott et al., 2023; Yanes Luis et al., 2023; Viseras and Garcia, 2019; Choi and Cielniak, 2021; Marchant et al., 2014; Denniston et al., 2023b) and occupancy grids (Yang et al., 2023b; Niroui et al., 2019; Chen et al., 2020; Lodel et al., 2022; Chen et al., 2019; Westheider et al., 2023; Liu et al., 2023; Gao and Zhang, 2022; Zwecher et al., 2022; Chen et al., 2022, 2021; Cao et al., 2023b; Sukkar et al., 2019), both of which are described in Sec. 3.1. Another common component of the state is the previous position and sensor history. The history components can range from the executed trajectory (Cao et al., 2023a) to the previous graph history (Cao et al., 2023c) as well as the previous observations (Viseras and Garcia, 2019; Denniston et al., 2023a). Some methods implicitly capture the history of robot poses and measurements in the belief state action-observation history (Choudhury et al., 2020; Ott et al., 2023) or by capturing the previously explored areas in the occupancy grid (Chen et al., 2021; Westheider et al., 2023; Gao and Zhang, 2022).
@@
Continuous action spaces often significantly increase the computational complexity of POMDPs. In many cases, discrete action spaces are simpler to work with. When the spatial map is represented as an abstract graph, a common choice is for movement commands to be represented as movement from one graph node to another via the connecting edge (Cao et al., 2023a, b, c; Yang et al., 2023a; Wei and Zheng, 2020; Rückin et al., 2022a; Ott et al., 2023; Choudhury et al., 2020; Westheider et al., 2023; Yanes Luis et al., 2023; Viseras and Garcia, 2019; Viseras et al., 2021; Bayerlein et al., 2021; Denniston et al., 2023a; Duecker et al., 2021; Denniston et al., 2023b; Gao and Zhang, 2022; Best et al., 2019; Zwecher et al., 2022). Another common approach relies on frontiers, which represent the boundary between known and unknown space in the environment and can be seen as a special case of a graph node action. The use of frontiers as movement actions does not necessarily require the spatial map to be represented as an abstract graph (Niroui et al., 2019; Chen et al., 2020, 2021; Liu et al., 2023). Other discrete actions include selecting from a discrete set of sensors to employ (Choudhury et al., 2020; Ott et al., 2023) as well as motion primitive commands (Marchant et al., 2014; Song et al., 2015) which can be seen as a discretization of continuous actions where the agent is given a set of potential trajectories to execute which correspond to actions.
@@
The reward function is the most problem-dependent component of the formulation. One of the core challenges that arise in the POMDP framework and, consequently, in AIPP, is the exploration-exploitation trade-off. The agent must select actions that balance between exploring unknown areas of the environment to increase its knowledge (exploration) and exploiting its current knowledge to find areas of interest within the constraints of its available resources (exploitation). This trade-off is intrinsically linked to the AIPP objective of maximizing the information criterion while respecting the cost constraint. Hence, it is a critical factor driving the learning process of the RL agent. Essentially, all AIPP methods use some measure of informativeness as the objective as discussed in Sec. 3.2. Map uncertainty reduction in high-interest areas is a very common RL reward in the AIPP literature (Yang et al., 2023a; Ott et al., 2023; Rückin et al., 2022a; Westheider et al., 2023; Viseras et al., 2021; Denniston et al., 2023a; Marchant et al., 2014; Duecker et al., 2021; Denniston et al., 2023b; Song et al., 2015). Some approaches use only map uncertainty reduction to spread measurements across the environment as uniformly as possible and minimize the estimation error Cao et al. (2023a); Niroui et al. (2019); Cao et al. (2023c); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Cao et al. (2023b). Coverage is a common surrogate metric for map uncertainty reduction since it only relies on computing the area uncovered as the robot travels through the environment, which is often faster than computing the Gaussian process variance online (Zeng et al., 2022; Liu et al., 2023; Kumar et al., 2022; Bayerlein et al., 2021; Sukkar et al., 2019; Best et al., 2019; Zwecher et al., 2022; Chen et al., 2022). While coverage can perform well on the map uncertainty reduction objective, it cannot explicitly reason about high-interest areas. The mutual information between observations and landmarks is another common way to measure map uncertainty reduction (Yang et al., 2023b; Lodel et al., 2022; Wei and Zheng, 2020; Chen et al., 2019). Some methods also explicitly incorporate distance-based metrics in the reward function (Chen et al., 2020, 2021; Cao et al., 2023b; Lodel et al., 2022; Gao and Zhang, 2022). For example, Chen et al. (2020, 2021) and Cao et al. (2023b) assign greater rewards for actions that achieve the same uncertainty reduction in less distance.
@@
Optimally solving POMDPs is computationally challenging (Kaelbling et al., 1998; Kochenderfer et al., 2022). However, practical solutions can be obtained through various RL algorithms. We categorize AIPP policies based on their deployment inference strategies, distinguishing between zero-shot inference and expected discounted $m$-step search policies. Zero-shot inference policies are trained offline, considering the sequence of actions. The policy is trained to account for the cascading impact of each action on subsequent decisions, taking into consideration the consequences of actions and future dependencies during the offline training process. Upon deployment, these policies take the current belief state—updated to reflect the cumulative knowledge of the environment as per Eq. 2—and directly output the next best action without additional planning. Conversely, policies using an expected discounted $m$-step search are dynamic in the sense that they actively consider a sequence of potential future actions during deployment. Instead of reacting to the current belief state, this approach projects forward, evaluating a decision tree of potential action sequences up to $m$ steps ahead and factoring in the anticipated updates to the belief state for each action. The policy then chooses an action based on the strategy that maximizes the total expected reward over this lookahead horizon.
@@
State
Current robot information
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Rückin et al. (2022a); Chen et al. (2019); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Sukkar et al. (2019); Best et al. (2019); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Arora et al. (2019); Bucher et al. (2021); Singh et al. (2009)
Observation history
Ye et al. (2018); Lodel et al. (2022); Cao et al. (2023c); Choudhury et al. (2020); Westheider et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Gazani et al. (2023); Denniston et al. (2023a); Bucher et al. (2021); Singh et al. (2009)
Robot history
Cao et al. (2023a); Chen et al. (2020, 2021); Cao et al. (2023c); Gazani et al. (2023); Denniston et al. (2023a); Gao and Zhang (2022)
Spatial map
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Rückin et al. (2022a); Chen et al. (2019); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Sukkar et al. (2019); Best et al. (2019); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Arora et al. (2019); Singh et al. (2009)
Map
2D Gaussian process
Cao et al. (2023a, c); Yang et al. (2023a); Wei and Zheng (2020); Rückin et al. (2022a); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Marchant et al. (2014); Denniston et al. (2023b)
2D occupancy grid
Yang et al. (2023b); Niroui et al. (2019); Chen et al. (2020); Lodel et al. (2022); Chen et al. (2019); Westheider et al. (2023); Liu et al. (2023); Gao and Zhang (2022); Zwecher et al. (2022); Chen et al. (2022)
3D occupancy grid
Chen et al. (2021); Cao et al. (2023b); Sukkar et al. (2019)
3D semantic occupancy map
(Zeng et al., 2022; Bartolomei et al., 2021)
Bayesian network
Arora et al. (2019)
Factor graph
Denniston et al. (2023a)
Gaussian Markov random field
Duecker et al. (2021)
Landmark locations and uncertainties
Chen et al. (2020, 2021); Bartolomei et al. (2021); Fan et al. (2020)
Location graph
Best et al. (2019); Choudhury et al. (2020); Hüttenrauch et al. (2019); Qie et al. (2019); Singh et al. (2009)
Action
Continuous action
Yang et al. (2023b); Lodel et al. (2022); Bartolomei et al. (2021); Sukkar et al. (2019); Hüttenrauch et al. (2019); Fan et al. (2020); Denniston et al. (2023b); Bucher et al. (2021)
Discrete action
Zeng et al. (2022); Chen et al. (2019); Ott et al. (2023); Viseras et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Marchant et al. (2014); Chen et al. (2022); Qie et al. (2019); Arora et al. (2019); Ye et al. (2018)
Frontier
Niroui et al. (2019); Chen et al. (2020, 2021); Liu et al. (2023)
Graph node
Cao et al. (2023a, b, c); Yang et al. (2023a); Wei and Zheng (2020); Rückin et al. (2022a); Ott et al. (2023); Choudhury et al. (2020); Westheider et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bayerlein et al. (2021); Denniston et al. (2023a); Duecker et al. (2021); Denniston et al. (2023b); Gao and Zhang (2022); Best et al. (2019); Zwecher et al. (2022); Singh et al. (2009)
Policy
Expected discounted $m$-step reward
Rückin et al. (2022a); Ott et al. (2023); Choudhury et al. (2020); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Sukkar et al. (2019); Best et al. (2019); Arora et al. (2019); Bucher et al. (2021)
Zero-shot inference
Yang et al. (2023b); Cao et al. (2023a); Niroui et al. (2019); Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Cao et al. (2023c); Yang et al. (2023a); Zeng et al. (2022); Wei and Zheng (2020); Chen et al. (2019); Westheider et al. (2023); Yanes Luis et al. (2023); Liu et al. (2023); Viseras and Garcia (2019); Viseras et al. (2021); Bartolomei et al. (2021); Kumar et al. (2022); Choi and Cielniak (2021); Bayerlein et al. (2021); Gazani et al. (2023); Gao and Zhang (2022); Hüttenrauch et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Qie et al. (2019); Fan et al. (2020); Denniston et al. (2023b); Ye et al. (2018); Singh et al. (2009)
Reward
Belief state mode
Choudhury et al. (2020)
Coverage
Zeng et al. (2022); Liu et al. (2023); Kumar et al. (2022); Bayerlein et al. (2021); Sukkar et al. (2019); Best et al. (2019); Zwecher et al. (2022); Chen et al. (2022); Bucher et al. (2021)
Maximize inter-agent distance
(Hüttenrauch et al., 2019)
Maximize mutual information of observations and landmarks
Yang et al. (2023b); Lodel et al. (2022); Wei and Zheng (2020); Chen et al. (2019)
Minimize landmark localization uncertainty
Chen et al. (2020, 2021); Bartolomei et al. (2021); Gazani et al. (2023)
Minimize map uncertainty
Cao et al. (2023a); Niroui et al. (2019); Cao et al. (2023c); Ott et al. (2023); Yanes Luis et al. (2023); Viseras and Garcia (2019); Choi and Cielniak (2021); Cao et al. (2023b)
Minimize map uncertainty in high-interest areas
Yang et al. (2023a); Ott et al. (2023); Rückin et al. (2022a); Westheider et al. (2023); Viseras et al. (2021); Denniston et al. (2023a); Marchant et al. (2014); Duecker et al. (2021); Denniston et al. (2023b); Arora et al. (2019); Singh et al. (2009)
Size of bounding box for target detection
Ye et al. (2018)
Travel distance
Chen et al. (2020, 2021); Cao et al. (2023b); Lodel et al. (2022); Gao and Zhang (2022); Qie et al. (2019)
Travel time
Fan et al. (2020)
@@
Yang et al. (2023b) use an attention-based multi-layer perceptron (MLP) which is trained using proximal policy optimization (PPO) to produce a zero-shot inference action of continuous longitudinal and lateral velocity with the goal of maximizing the mutual information between observations and landmark states. Cao et al. (2023a) take a slightly different approach by using an attention-based graph neural network and long short-term memory (LSTM) network with PPO to select the next graph node. Rückin et al. (2022a) use AlphaZero by combining Monte Carlo tree search (MCTS) with a convolutional neural network (CNN) to learn information-rich actions in adaptive data gathering missions. Chen et al. (2019) uses deep Q-network (DQN) with a CNN to select the next location to visit from a set of randomly sampled 2D positions with the goal of maximizing the mutual information between received observations and the occupancy grid map. Chen et al. (2021) uses the advantage actor-critic (A2C) algorithm with a graph neural network to select the next best frontier to visit with the goal of maximizing landmark uncertainty reduction and minimizing travel distance. Cao et al. (2023b) uses the soft actor-critic (SAC) algorithm with an attention-based graph neural network to output the next graph node to visit with the goal of maximizing the number of observed frontiers while minimizing travel distance. Other methods mainly differ in the training algorithms and the network architectures used with the most popular being PPO, A2C, SAC, and CNN, DQN, LSTM, respectively (Yang et al., 2023b; Cao et al., 2023a; Niroui et al., 2019; Chen et al., 2020, 2021; Cao et al., 2023b; Lodel et al., 2022; Cao et al., 2023c; Chen et al., 2019; Liu et al., 2023; Gazani et al., 2023; Zwecher et al., 2022).

### v3

@@ 4.2 Reinforcement Learning
The problem of AIPP can be framed as a partially observable Markov decision process (POMDP)  [112]. Formally, a POMDP is defined by the tuple $\langle\mathcal{S},\mathcal{A},\mathcal{O},T,Z,R\rangle$, where:
@@
In the context of AIPP, the true state of the environment is the actual physical attributes of the environment captured by $F(\mathbf{x})$ in Sec. 2, e.g., the locations of obstacles and areas of interest, while the observation $z$ is the robot’s current measurement of these attributes based on sensor readings.
@@
The state transition function $T(s^{\prime}\,|\,s,a)$ describes the probabilistic dynamics of the robot’s state in the context of the AIPP problem. When the robot executes an action $a\in\mathcal{A}$ in its current state $s\in\mathcal{S}$, it moves to a new state $s^{\prime}\in\mathcal{S}$. The inherent uncertainty of action outcomes and their impact on the environment is captured by the transition function. This function is used in the belief update, which refines the robot’s understanding of the environment feature mapping $F$. Additionally, the observation function $Z(o\,|\,s^{\prime},a)$ provides the likelihood of perceiving an observation $o\in\mathcal{O}$ subsequent to action $a$ and arriving at state $s^{\prime}$. This observation model, which is probabilistically linked to the feature mapping $F$ through $z_{t}\sim p(z\,|\,F)$, is used to update the robot’s belief state using Eq. 2.
@@
Any POMDP can be viewed as an MDP that uses beliefs as states, also called a belief-state MDP or belief MDP. The state space of a belief MDP is the set of all beliefs $\mathcal{B}$. The action space is equivalent to that of the POMDP [112]. To connect this directly to Sec. 3, in AIPP, the robot’s belief about the true state of the environment is denoted by $\tilde{F}$ in Sec. 3.1. This belief evolves as the robot gathers information about its environment, which starts as partially unknown and is gradually discovered through sensing and exploration.
4.2.1 Network Architectures and Training Algorithms
Various neural network architectures can be used in RL for AIPP, such as attention-based neural networks, convolutional neural networks (CNN), long short-term memory (LSTM) networks, and graph neural networks (GNN).
Attention-based neural networks, inspired by human visual attention, focus selectively on input data, which is useful for AIPP where the robot focuses on high-interest regions. CNNs excel at capturing spatial hierarchies in 2D occupancy grid maps. LSTMs, a type of recurrent neural network, learn long-term dependencies, which is ideal for temporally dependent problems in AIPP. GNNs operate on graph structures, making them most appropriate for graph-modeled environments.
Several RL algorithms can train these network architectures including proximal policy optimization (PPO), asynchronous actor-critic (A3C), soft actor-critic (SAC), advantage actor-critic (A2C), deep Q-network (DQN), double deep Q-network (DDQN), as well as many other variations and combinations, e.g. AlphaZero. These algorithms vary in exploration and exploitation, sample efficiency, computational needs, and stability.
Actor-critic methods use a value function estimate to guide optimization. The actor is the policy, while the critic is the value function. They train in parallel, differing in value function, advantage function, or action-value function approximation. Most methods focus on stochastic policies, but some support deterministic continuous actions. For example, PPO, A3C, SAC, and A2C follow an actor-critic approach with separate policy and value networks. In contrast, DQN and DDQN are critic-only value-based methods, where the optimal policy is inferred from the value function.
AlphaZero combines the actor-critic framework with Monte Carlo tree search (MCTS), using the network’s policy to guide the search and the value function to evaluate leaf nodes. MCTS simulations generate a robust policy, considering a longer planning horizon [113, 114].
Various open-source implementations and benchmarks of these network architectures and training algorithms are widely available  [115, 116, 117, 118, 119].
4.2.2 Reinforcement Learning Approaches
@@
All of the methods surveyed include some aspect of the current robot information in the state. The robot pose is often used [41, 32, 43, 71, 59], while other methods include additional robot information, e.g., energy level [7, 62], viewing direction [50], and operational status [120]. In addition to current robot information, all methods store some form of a spatial map in the state. The most popular spatial maps are Gaussian processes [2, 70, 63, 66, 5, 62, 64, 6, 69, 71, 59] and occupancy grids [121, 40, 29, 32, 33, 27, 46, 53, 54, 50, 34, 41, 30], both of which are described in Sec. 3.1. Another common component of the state is the previous position and sensor history. The history components can range from the executed trajectory [2] to the previous graph history [70] as well as the previous observations [6, 4]. Some methods implicitly capture the history of robot poses and measurements in the belief state action-observation history [7, 62] or by capturing the previously explored areas in the occupancy grid [34, 27, 53].
@@
Continuous action spaces often significantly increase the computational complexity of POMDPs. In many cases, discrete action spaces are simpler to work with. When the spatial map is represented as an abstract graph, a common choice is for movement commands to be represented as movement from one graph node to another via the connecting edge [2, 41, 70, 63, 66, 5, 62, 7, 27, 64, 6, 35, 120, 4, 73, 59, 53, 75, 54]. Another common approach relies on frontiers, which represent the boundary between known and unknown space in the environment and can be seen as a special case of a graph node action. The use of frontiers as movement actions does not necessarily require the spatial map to be represented as an abstract graph [40, 29, 34, 46]. Other discrete actions include selecting from a discrete set of sensors to employ [7, 62] as well as motion primitive commands [71, 60] which can be seen as a discretization of continuous actions where the agent is given a set of potential trajectories to execute which correspond to actions.
@@
The reward function is the most problem-dependent component of the formulation. One of the core challenges that arise in the POMDP framework and, consequently, in AIPP, is the exploration-exploitation trade-off. The agent must select actions that balance between exploring unknown areas of the environment to increase its knowledge (exploration) and exploiting its current knowledge to find areas of interest within the constraints of its available resources (exploitation). This trade-off is intrinsically linked to the AIPP objective of maximizing the information criterion while respecting the cost constraint. Hence, it is a critical factor driving the learning process of the RL agent. Essentially, all AIPP methods use some measure of informativeness as the objective as discussed in Sec. 3.2. Map uncertainty reduction in high-interest areas is a very common RL reward in the AIPP literature [63, 62, 5, 27, 35, 4, 71, 73, 59, 60]. Some approaches use only map uncertainty reduction to spread measurements across the environment as uniformly as possible and minimize the estimation error [2, 40, 70, 62, 64, 6, 69, 41]. Coverage is a common surrogate metric for map uncertainty reduction since it only relies on computing the area uncovered as the robot travels through the environment, which is often faster than computing the Gaussian process variance online [43, 46, 98, 120, 30, 75, 54, 50]. While coverage can perform well on the map uncertainty reduction objective, it cannot explicitly reason about high-interest areas. The mutual information between observations and landmarks is another common way to measure map uncertainty reduction [121, 32, 66, 33]. Some methods also explicitly incorporate distance-based metrics in the reward function [29, 34, 41, 32, 53]. For example, Chen et al. [29, 34] and Cao et al. [41] assign greater rewards for actions that achieve the same uncertainty reduction in less distance.
@@
Optimally solving POMDPs is computationally challenging [123, 112]. However, practical solutions can be obtained through various RL algorithms. We categorize AIPP policies based on their deployment inference strategies, distinguishing between zero-shot inference and expected discounted $m$-step search policies. Zero-shot inference policies are trained offline, considering the sequence of actions. The policy is trained to account for the cascading impact of each action on subsequent decisions, taking into consideration the consequences of actions and future dependencies during the offline training process. Upon deployment, these policies take the current belief state—updated to reflect the cumulative knowledge of the environment as per Eq. 2—and directly output the next best action without additional planning. Conversely, policies using an expected discounted $m$-step search are dynamic in the sense that they actively consider a sequence of potential future actions during deployment. Instead of reacting to the current belief state, this approach projects forward, evaluating a decision tree of potential action sequences up to $m$ steps ahead and factoring in the anticipated updates to the belief state for each action. The policy then chooses an action based on the strategy that maximizes the total expected reward over this lookahead horizon.
@@
State Current robot information [121, 2, 40, 29, 34, 41, 32, 70, 63, 43, 66, 5, 33, 62, 7, 27, 64, 46, 6, 35, 122, 98, 69, 120, 93, 4, 71, 73, 59, 53, 30, 75, 72, 54, 50, 124, 125, 126, 127, 8, 128, 129]
Observation history [130, 32, 70, 7, 27, 46, 6, 93, 4, 127, 8]
Robot history [2, 29, 34, 70, 93, 4, 53]
Spatial map [121, 2, 40, 29, 34, 41, 32, 70, 63, 43, 66, 5, 33, 62, 7, 27, 64, 46, 6, 35, 122, 98, 69, 120, 93, 4, 71, 73, 59, 53, 30, 75, 72, 54, 50, 124, 125, 126, 8, 128, 129]
Map 2D Gaussian process [2, 70, 63, 66, 5, 62, 64, 6, 69, 71, 59]
2D occupancy grid [121, 40, 29, 32, 33, 27, 46, 53, 54, 50, 128, 129]
3D occupancy grid [34, 41, 30]
3D semantic occupancy map [43, 122]
Bayesian network [126]
Factor graph [4]
Gaussian Markov random field [73]
Landmark locations and uncertainties [29, 34, 122, 125]
Location graph [75, 7, 72, 124, 8, 128, 129]
Action Continuous action [121, 32, 122, 30, 72, 125, 59, 127]
Discrete action [43, 33, 62, 35, 98, 69, 120, 93, 71, 50, 124, 126, 130, 128]
Frontier [40, 29, 34, 46]
Graph node [2, 41, 70, 63, 66, 5, 62, 7, 27, 64, 6, 35, 120, 4, 73, 59, 53, 75, 54, 8, 129]
Policy Expected discounted $m$-step reward [5, 62, 7, 4, 71, 73, 59, 30, 75, 126, 127, 128, 129]
Zero-shot inference [121, 2, 40, 29, 34, 41, 32, 70, 63, 43, 66, 33, 27, 64, 46, 6, 35, 122, 98, 69, 120, 93, 53, 72, 54, 50, 124, 125, 59, 130, 8]
Reward Belief state mode [7]
Coverage [43, 46, 98, 120, 30, 75, 54, 50, 127, 128, 129]
Maximize inter-agent distance [72]
Maximize mutual information of observations and landmarks [121, 32, 66, 33]
Minimize landmark localization uncertainty [29, 34, 122, 93]
Minimize map uncertainty [2, 40, 70, 62, 64, 6, 69, 41]
Minimize map uncertainty in high-interest areas [63, 62, 5, 27, 35, 4, 71, 73, 59, 126, 8]
Size of bounding box for target detection [130]
Travel distance [29, 34, 41, 32, 53, 124]
Travel time [125]
@@
Yang et al. [121] use an attention-based multi-layer perceptron (MLP) which is trained using proximal policy optimization (PPO) to produce a zero-shot inference action of continuous longitudinal and lateral velocity with the goal of maximizing the mutual information between observations and landmark states. Cao et al. [2] take a slightly different approach by using an attention-based graph neural network and long short-term memory (LSTM) network with PPO to select the next graph node. Rückin et al. [5] use AlphaZero by combining Monte Carlo tree search (MCTS) with a convolutional neural network (CNN) to learn information-rich actions in adaptive data gathering missions. Chen et al. [33] uses deep Q-network (DQN) with a CNN to select the next location to visit from a set of randomly sampled 2D positions with the goal of maximizing the mutual information between received observations and the occupancy grid map. Chen et al. [34] uses the advantage actor-critic (A2C) algorithm with a graph neural network to select the next best frontier to visit with the goal of maximizing landmark uncertainty reduction and minimizing travel distance. Cao et al. [41] uses the soft actor-critic (SAC) algorithm with an attention-based graph neural network to output the next graph node to visit with the goal of maximizing the number of observed frontiers while minimizing travel distance. Other methods mainly differ in the training algorithms and the network architectures used with the most popular being PPO, A2C, SAC, and CNN, DQN, LSTM, respectively [121, 2, 40, 29, 34, 41, 32, 70, 33, 46, 93, 54].